# CIE: Controlling Language Model Text Generations Using Continuous Signals

### v1

@@ Abstract
Aligning language models with user intent is becoming increasingly relevant to enhance user experience. This calls for designing methods that can allow users to control the properties of the language that LMs generate. For example, controlling the length of the generation, the complexity of the language that gets chosen, the sentiment, tone, etc. Most existing work attempts to integrate users’ control by conditioning LM generations on natural language prompts or discrete control signals, which are often brittle and hard to scale. In this work, we are interested in continuous control signals, ones that exist along a spectrum that can’t easily be captured in a natural language prompt or via existing techniques in conditional generation. Through a case study in controlling the precise response-length of generations produced by LMs, we demonstrate how after fine-tuning, behaviors of language models can be controlled via continuous signals—as vectors that are interpolated between a “low” and a “high” token embedding. Our method more reliably exerts response-length control than in-context learning methods or fine-tuning methods that represent the control signal as a discrete signal. Our full open-sourced code and datasets are available at https://github.com/vsamuel2003/CIE.  [^1] Correspondence: vsamuel@umd.edu
@@
Vinay Samuel1,2  Harshita Diddee2  Yiming Zhang2  Daphne Ippolito2 1University of Maryland, College Park, 2Carnegie Mellon University
@@ 1 Introduction
Instruction-tuned language models have demonstrated remarkable capabilities in generating coherent responses to a given user instruction (Taori et al., 2023; Ouyang et al., 2022; Shu et al., 2024). However, users often want to influence specific properties of the generated text beyond just the content—for example, controlling the length of the generation (as illustrated in Figure 1), the complexity of the language, the sentiment, or the tone.
Figure 1: We present an example of CIE for controlling response-length. The user inputs both an instruction as well as a desired answer length. The control embedding for this response length is interpolated between the lower and upper bound control embeddings which are learned during training. Finally the fully augmented model embeddings are used to generate an attribute controlled response.
There are many approaches to controlling such attributes in language model outputs. While discrete control signals (such as special tokens or natural language descriptions) have shown promise (Konen et al., 2024; Deng et al., 2022; Keskar et al., 2019; Dathathri et al., 2019), they have inherent limitations. Discrete approaches struggle to provide fine-grained control over generated text properties or require extensive training to achieve competence (Keskar et al., 2019; Li et al., 2024). Meanwhile, natural language prompt templates tend to be brittle, with small changes potentially leading to inconsistent results (Zhuo et al., 2024).
@@
While similar approaches have demonstrated effectiveness in domains such as strength-conditioning in chess (Zhang et al., 2025), we extend this methodology to precisely control specified attributes in language model outputs. We utilize embeddings for incorporating continuous control signals rather than natural language instructions due to LMs’ limited numerical understanding (Yang et al., 2025). This approach offers advantages over discrete control methods by allowing smoother transitions between attribute intensities and more precise steering of generative outputs.
The core mechanism of CIE is a control embedding that is appended to the token embeddings of the user instruction. We augment the LM’s embedding matrix with two embeddings corresponding to the lowest and highest possible values for a given attribute. During fine-tuning and inference, the control embedding is computed through linear interpolation between these low and high embeddings. Given a dataset of instruction-answer pairs annotated with control values, CIE enables LMs to learn a mapping from the control embedding’s position to the attribute’s degree of control.
Our main contribution is CIE, a continuous control embedding approach to steering generation. To demonstrate its effectiveness, we apply CIE to response length control (as measured by word count) and compare our results with both a prompting baseline and a state-of-the-art discrete signal approach (Li et al., 2024).  [^2] We also open-source VerbosityCTRL which is used to demonstrate CIE’s efficacy for response length control.
@@ 2 Method
where $\mathbf{e}_{\text{lower}}$ represents the embedding vector for the lower bound of control values for $a$ in the training data $c_{\text{lower}}$ and $\mathbf{e}_{\text{upper}}$ is the embedding vector for the upper bound of the allowed values of $a$ in the training data $c_{\text{upper}}$. For a given $c$, we calculate the control embedding vector for $c$ as $\mathbf{e}_{\text{c}}=\alpha\mathbf{e}_{\text{lower}}+(1-\alpha)\mathbf{e}_{\text{upper}}$ where $\alpha=\frac{c_{\text{upper}}-c}{c_{\text{upper}-c_{\text{lower}}}}$. At training time, this control embedding vector $\mathbf{e}_{\text{c}}$ is appended to the embeddings for the user instruction in the instruct/chat templates of the models.
The training dataset $\mathcal{D}$ comprises triplets of $(i,a,wc)$, where $i$ represents the instruction, $a$ denotes the ground truth answer, and $wc$ indicates the conditioned response length. During training, $wc$ is set to match the response length in $a$. We establish engineering-determined lower and upper bounds ($c_{\text{lower}}$ and $c_{\text{upper}}$) for the controlled attribute, with all $wc$ values clamped within this range before calculating $e_{c}$. Following the approach of Zhang et al. (2025)., we curate $\mathcal{D}$ to maintain an approximately uniform distribution of response lengths, ensuring sufficient training for $\mathbf{e}_{\text{lower}}$ and $\mathbf{e}_{\text{upper}}$.
@@ 3.2 Metrics
We evaluate our results using the Conditioning Precision Ratio (CPR) metric, which measures how accurately models follow response length instructions. CPR simply checks for an exact match between the generated response length and the conditioned response length, returning 1 for a perfect match and 0 otherwise. We also define CPR @k, which introduces a relative tolerance factor, accepting response lengths within ±$\pm k$ percent of the target (e.g., CPR @0.1 allows deviations within 10% of the specified response length). To verify that CIE preserves language generation quality, we calculate win-rates between model outputs and a prompt baseline, counting both wins and ties as evidence that CIE does not harm generation abilities. Additional details on this evaluation approach are available in Appendix H.
@@ 4 Results
Our CIE approach significantly improved response length control across all evaluated models and datasets. On VerbosityCTRL ${}_{\text{val}}$, CIE increased average CPR from 0.80-9.80% to 8.90-9.80%, with substantial gains for initially weaker models like gemma-7B-IT (0.80% to 8.90%). The most dramatic improvements occurred at the CPR @0.05 threshold, with increases of 23.10, 24.00, and 30.40 percentage points for LLaMA-3-8B-IT, gemma-7B-IT, and Qwen-1.5-7B respectively. Similar patterns emerged on VerbosityCTRL ${}_{\text{range}}$, with average CPR rising from 1.6% to 4.8%. These improvements generalized to out-of-distribution data (Alpaca-LI), confirming CIE’s cross-contextual robustness. We observed an inverse relationship between baseline capability and improvement magnitude—weaker models achieved more dramatic relative gains, with gemma-7B-IT showing over 10$\times$ improvement at exact-match CPR. Stronger models like LLaMA-3-8B-IT made minor sacrifices at exact-match for substantial gains at relaxed thresholds. The diminishing relative gains at CPR @0.1 suggest base models already capture many approximate matches at looser tolerances. Detailed analyses are available in Appendices F and G.
@@
CIE consistently outperforms Ruler across all model variants. For gemma-7B-IT, CIE achieves absolute improvements of 9.09 and 7.86 points in PM and FM scores, respectively. LLaMA-3-8B shows even more substantial gains of 16.30 and 16.18 points, while Qwen1.5-7B demonstrates improvements of 6.54 and 6.44 points. These significant performance gains highlight CIE’s effectiveness in enhancing precision response length conditioning, suggesting it provides a more robust framework for output length control compared to state-of-the-art methods.
@@ 6 Conclusion
Fine-grained control over language model outputs represents a critical capability for deploying these systems in contexts requiring adaptable text generation. While the vast majority of current controllability approaches are based on discrete signals, we believe that the fine-grained controllability of continuous signals is necessary for the evolving user demands of conversational LMs. Through our proposed method, we have established a framework that effectively modulates response length while preserving content fidelity. The success of this approach suggests promising applications beyond length control, particularly for other continuous or ordinal attributes such as linguistic complexity, descriptive density, and even subjective qualities like humor intensity or emotional resonance. Future work should explore how these continuous embedding controls generalize across diverse attributes and investigate potential compositionality when multiple control dimensions are adjusted simultaneously. Our findings indicate that embedding-based continuous signal controls may provide a more robust foundation for the next generation of highly adaptable language generation systems.
@@ Acknowledgments
The authors acknowledge AI assistance as follows: Several sections underwent refinement using AI, where the AI was instructed to ’rewrite in clear, coherent, and concise academic style writing while not altering the major points in the provided writing.’ All AI-generated content was thoroughly reviewed by the authors before inclusion in the manuscript, ensuring full compliance with ACL ARR guidelines. Additional some figure components (such as the model and user avatars) were either AI generated or AI refined. We also thank Barry Wang (CMU LTI PhD) for major inputs in designing the teaser figure.
@@ Appendix C Dataset Statistics
Table 5: Word-count statistics for the datasets used in our experiments.
@@ D.1 Hyperparameter Search
We perform a hyperparameter grid search for each tested model for the epochs and learning_rate hyperparameters. We created a random subset of VerbosityCTRL ${}_{\text{train}}$ with 10000 samples that was used for the grid search. We searched over epoch values of $[3,5,7]$ and learning_rate values of $[5\times 10^{-6},1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}]$ and evaluated performance on VerbosityCTRL ${}_{\text{val}}$ to determine the epoch and learning_rate for each model.
@@ Appendix F Detailed Results Analysis
We observed an inverse relationship between baseline capability and relative improvement with CIE. LLaMA-3-8B-IT demonstrated the highest baseline control on VerbosityCTRL ${}_{\text{val}}$ (9.80 CPR), followed by Qwen-1.5-7B (4.60 CPR) and Gemma-7B-IT (0.80 CPR). However, initially weaker models achieved the most dramatic gains: Gemma-7B-IT showed over 10$\times$ improvement at CPR (0.80 to 8.90) while LLaMA-3-8B-IT made slight sacrifices at exact-match for substantial gains at relaxed thresholds. These improvements generalized across both VerbosityCTRL ${}_{\text{range}}$ and out-of-distribution Alpaca-LI evaluation sets, confirming the method’s robustness and effectiveness with initially weaker models.

### v2

@@ Abstract
Aligning language models (LMs) with user intent is becoming increasingly relevant to enhance user experience. This calls for designing methods that can allow users to control the properties of the language that LMs generate, for example, controlling the length of the generation or the complexity of the language that gets chosen. Most existing work attempts to integrate users’ control by conditioning LM generations on natural language prompts or discrete control signals, which are often brittle and hard to scale. In this work, we are interested in continuous control signals, ones that exist along a spectrum that can’t easily be captured in a natural language prompt or via existing techniques in conditional generation. Through a case study in controlling the precise response-length of generations, we demonstrate how an LM can be finetuned to expect a control vector that is interpolated between a “low” and a “high” token embedding. Our method more reliably exerts response-length control than in-context learning methods or fine-tuning methods that represent the control signal as a discrete signal.
@@
Vinay Samuel1,2,**footnotemark: * Harshita Diddee2Yiming Zhang2Daphne Ippolito2 1University of Maryland, College Park, 2Carnegie Mellon University
00footnotetext: * Correspond to vsamuel@umd.edu.
@@ 1 Introduction
Instruction-tuned language models have demonstrated remarkable capabilities in generating coherent responses to user instructions (Taori et al., 2023; Ouyang et al., 2022; Shu et al., 2024). However, users often want to influence specific properties of the generated text beyond just the content–for example, controlling the length of the generation (as illustrated in Figure 1), the complexity of the language, the sentiment, or the tone.
Figure 1: When using CIE for controlling response length, the user inputs both an instruction as well as a desired answer length. A control embedding for this response length is interpolated between the lower and upper bound control embeddings which were learned during training. The control embedding is appended to the input token embedding sequence of the instruction, and the LM which has been finetuned to expect this input generates an attribute-controlled response.
There are many approaches to controlling such attributes in language model outputs. While discrete control signals (such as special tokens or words that get added to the user’s prompt) have shown promise (Konen et al., 2024; Deng et al., 2022; Keskar et al., 2019; Dathathri et al., 2019), they have inherent limitations. Discrete approaches struggle to provide fine-grained control (especially for properties that exist on a continuous spectrum), and they can require extensive training to achieve competence (Keskar et al., 2019; Li et al., 2024). Meanwhile, methods that achieve controllability through natural-language instructions tend to be brittle, with small, immaterial changes to the verbalization potentially leading to inconsistent results (Zhuo et al., 2024).
@@
While similar approaches have demonstrated effectiveness in domains such as strength-conditioning in chess (Zhang et al., 2025), we extend this methodology to precisely control specified attributes in language model outputs. Using continuous embeddings for incorporating continuous control signals rather than natural language instructions allows us to circumvent LLMs’ limited numerical understanding (Yang et al., 2025). Our approach also offers advantages over discrete control token methods by allowing smoother transitions between attribute intensities and more precise steering of generative outputs.
The core mechanism of CIE is a control embedding that is appended to the token embeddings of the user instruction. We augment the LM’s existing embedding matrix with two new embeddings corresponding to the lowest and highest possible values for a given attribute. During fine-tuning and inference, the control embedding is computed through linear interpolation between these low and high embeddings. Given a dataset of instruction-answer pairs annotated with control values, CIE enables LMs to learn a mapping from the control embedding’s position to the attribute’s degree of control.
To demonstrate our proposed method CIE’s effectiveness, we apply it to response length control (as measured by word count) and compare our results with both a prompting baseline and a state-of-the-art discrete signal approach (Li et al., 2024).  [^2] Our open-source code and training data are available at https://github.com/vsamuel2003/CIE.
@@ 2 Method
where $\mathbf{e}_{\text{lower}}$ represents the embedding vector for the lower bound of control values for $a$ in the training data $c_{\text{lower}}$ and $\mathbf{e}_{\text{upper}}$ is the embedding vector for the upper bound of the allowed values of $a$ in the training data $c_{\text{upper}}$. For a given $c$, we calculate the control embedding vector for $c$ as $\mathbf{e}_{\text{c}}=\alpha\mathbf{e}_{\text{lower}}+(1-\alpha)\mathbf{e}_{\text{upper}}$ where $\alpha=\frac{c_{\text{upper}}-c}{c_{\text{upper}}-c_{\text{lower}}}$.
A <control-embedding> token is appended to instructions without expanding vocabulary. During the forward pass, control positions are replaced with interpolated embeddings computed from two learned embeddings ($\mathbf{e}_{\text{lower}}$ and $\mathbf{e}_{\text{upper}}$) before transformer processing. The interpolated control embedding provides a continuous conditioning signal that influences all subsequent transformer layers. Unlike discrete tokens that compete for attention with content tokens, the control embedding acts as a persistent bias that guides generation decisions throughout the decoding process.
The control embedding is interpolated between trained bounds and injected into instruction embeddings for standard autoregressive generation. During decoding, the control embedding establishes a representation space bias that influences next-token probability distributions. The learned embeddings encode attribute-specific patterns that, when interpolated, provide fine-grained control over generation characteristics.
The training dataset $\mathcal{D}$ comprises triplets of $(i,a,wc)$, where $i$ represents the instruction, $a$ denotes the ground truth answer, and $wc$ indicates the conditioned response length. During training, $wc$ is set to match the response length in $a$. We establish engineering-determined lower and upper bounds ($c_{\text{lower}}$ and $c_{\text{upper}}$) for the controlled attribute, with all $wc$ values clamped within this range before calculating $e_{c}$. Following the approach of Zhang et al. (2025), we curate $\mathcal{D}$ to maintain an approximately uniform distribution of response lengths, ensuring sufficient training for $\mathbf{e}_{\text{lower}}$ and $\mathbf{e}_{\text{upper}}$.
The method adds only a $(2\times D)$ control embedding matrix to the base model, requiring minimal modifications to standard causal language modeling.
@@ 3.2 Metrics
We evaluate our results using the Conditioning Precision Ratio (CPR) metric, which measures how accurately models follow response length instructions. CPR simply checks for an exact match between the generated response length and the conditioned response length, returning 1 for a perfect match and 0 otherwise. We also define CPR @k, which introduces a relative tolerance factor, accepting response lengths within $\pm k$ percent of the target (e.g., CPR @0.1 allows deviations within 10% of the specified response length). To verify that CIE preserves language generation quality, we calculate win-rates between model outputs and a prompt baseline, counting both wins and ties as evidence that CIE does not harm generation abilities. Additional details on this evaluation approach are available in Appendix H.
@@ 4 Results
9.80
9.50 $\downarrow\,0.30$
22.70
45.80 $\uparrow\,23.10$
38.90
72.70 $\uparrow\,33.80$
3.19
5.05 $\uparrow\,1.86$
17.40
39.08 $\uparrow\,21.68$
39.04
71.65 $\uparrow\,32.61$
5.43
9.04 $\uparrow\,3.61$
21.95
45.02 $\uparrow\,23.07$
43.67
72.40 $\uparrow\,28.73$
gemma-7B-IT
@@
0.80
8.90 $\uparrow\,8.10$
4.90
28.90 $\uparrow\,24.00$
8.80
47.50 $\uparrow\,38.70$
0.73
4.42 $\uparrow\,3.69$
3.87
37.90 $\uparrow\,34.03$
9.94
68.37 $\uparrow\,58.43$
1.58
4.75 $\uparrow\,3.17$
9.05
40.95 $\uparrow\,31.90$
17.19
65.16 $\uparrow\,47.97$
Qwen-1.5-7B
@@
4.60
9.80 $\uparrow\,5.20$
9.40
39.80 $\uparrow\,30.40$
16.60
67.80 $\uparrow\,51.20$
0.91
4.77 $\uparrow\,3.86$
5.96
31.55 $\uparrow\,25.59$
14.82
62.71 $\uparrow\,47.89$
2.94
4.20 $\uparrow\,1.26$
8.82
20.36 $\uparrow\,11.54$
17.42
38.46 $\uparrow\,21.04$
@@
[…]
@@
CIE consistently outperforms Ruler across all model variants. For gemma-7B-IT, CIE achieves absolute improvements of 9.09 and 7.86 points in PM and FM scores, respectively. LLaMA-3-8B shows even more substantial gains of 16.30 and 16.18 points, while Qwen1.5-7B demonstrates improvements of 6.54 and 6.44 points.
To understand why CIE outperforms discrete methods, we conducted scaling experiments training LLaMA-3-8B on both RULER and CIE at 25%, 50%, 75%, and 100% of training data. Results from Table 2 show CIE maintains superiority across all data scales without convergence. CIE’s continuous interpolation creates a structured embedding space where similar lengths have similar representations, enabling generalization between target values with fewer examples per length. The continuous space inherently captures ordinal relationships between control values, whereas discrete methods lack this geometric structure. Additionally, the requirement of adding in additional MLT tokens in the RULER method makes the method non-trivial to scale to larger word counts.
These significant performance gains highlight CIE’s effectiveness in enhancing precision response length conditioning, suggesting it provides a more robust framework for output length control compared to state-of-the-art methods.
Model 25% 50% 75% 100%
LLaMA-3-8B_Ruler 11.80/14.20 24.20/28.00 38.00/42.50 49.22/53.33
LLaMA-3-8B_CIE 13.08/16.32 26.72/30.38 41.46/44.79 65.52/69.51
Table 2: Performance Analysis: Why CIE Outperforms Discrete Methods. The results show Precise Math / Flexible Match
@@ 6 Conclusion
Fine-grained control over language model outputs represents a critical capability for deploying these systems in contexts requiring adaptable text generation. While the vast majority of current controllability approaches are based on discrete signals, we believe that the fine-grained controllability of continuous signals is necessary for the evolving user demands of conversational LMs. Through our proposed method, we have established a framework that effectively modulates response length while preserving content fidelity.
Beyond word count, future work should aim to apply CIE as well as other continuous signal approaches to control concrete attributes like sentence count, character count, and complexity measures. While subjective properties like toxicity and bias present additional challenges in defining meaningful continuous ranges, as discussed in the Limitations of this paper, they represent important directions for future exploration.
@@ Acknowledgments
The authors acknowledge AI assistance as follows: Several sections underwent refinement using AI, where the AI was instructed to ’rewrite in clear, coherent, and concise academic style writing while not altering the major points in the provided writing.’ All AI-generated content was thoroughly reviewed by the authors before inclusion in the manuscript, ensuring full compliance with ACL ARR guidelines. Additionally some figure components (such as the model and user avatars) were either AI generated or AI refined. We also thank Barry Wang (CMU CSD PhD) for major inputs in designing the teaser figure.
@@ Appendix C Dataset Statistics
Table 6: Word‐count statistics for the datasets used in our experiments.
@@ D.1 Hyperparameter Search
We perform a hyperparameter grid search for each tested model for the epochs and learning_rate hyperparameters. We created a random subset of VerbosityCTRL $_{\text{train}}$ with 10000 samples that was used for the grid search. We searched over epoch values of $[3,5,7]$ and learning_rate values of $5\times 10^{-6},1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}]$ and evaluated performance on VerbosityCTRL $_{\text{val}}$ to determine the epoch and learning_rate for each model.
@@ Appendix F Detailed Results Analysis
[…]